# Data for Mathematical Copilots: Better Ways of Presenting Proofs for Machine Learning

### v1

@@ 4.3 Motivated Proofs
We do not doubt that there is substantial value in bespoke AI tools that can automate specific elements of a mathematician’s workflow. However, if the goal is not only to provide correct proofs but also to enhance a user’s mathematical understanding of the results (Zhang et al., 2023a) and guide new discoveries, we need new tooling – and we argue, new kinds of data. In particular, we argue for data which faithfully represents the process of proof discovery instead of proof exposition. Towards this goal, we introduce motivated proofs (Pólya, 1949; Morris, 2019), which contain and make transparent more of the proof discovery process. We argue for their value as a standard for LLM evaluation. We then provide some examples and discuss some preliminary observations about current LLM’s ability to construct and identify motivated proofs. Finally, we speculate on a path towards large-scale evaluation of a model’s ability to produce motivated proofs.
@@ 4.3.1 Examples of motivated proofs
But we also know that $x$ is nilpotent, so if we have $r$ such that $x^{r}=0$, then all coefficients from $x^{r}$ onwards can be ignored. This solves our issue, as we can let $m=r-1$ to remove the $a_{m}=0$ condition, and we are left with $a_{k}=(-1)^{k}$, so $(1+x)$ is a unit with inverse $\sum_{k=0}^{r-1}(-1)^{k}\cdot x^{k}$.
@@ 4.3.2 Motivated proofs as an evaluation metric
As argued in  (Pólya, 1949; Morris, 2019), a motivated proof is more informative to the reader, as it provides more insight into how the proof was discovered and how the reader might discover similar results themselves. More speculatively, a model capable of producing motivated proofs will likely generalize better to novel problems. The first reason for this conjecture is that most results in the literature are not fully motivated, and a model that is capable of recognising this is less likely to blindly recall its training data. Requiring proofs to be motivated also restricts the amount of brute-force search a model can perform, so performance depends more on reasoning ability and less on test-time computational power.
@@
In most responses, the model showed some evidence of “understanding” the concept of motivated proofs by avoiding unmotivated constructions. However, this often led to pseudo-motivations, where some justification was given before a construction, but it didn’t provide any useful substance besides obscuring the unmotivated step (and often appealed to some irrelevant fallacies). In the example below, the unmotivated step of Cantor’s theorem is relegated to the statement “An important question arises: Does $x$ belong to its own image $f(x)$?”.
@@
Overall, o1-preview seemed to outperform Llama on most problems. We also observed a slight improvement by using negative as well as positive examples to “explain” the concept. As an unfortunate side effect, the extra information also increased the tendency to “obscure” unmotivated arguments as discussed above.
@@
To conduct a larger scale investigation of a model’s ability to generate motivated proofs, it would help if there was an existing “judge” model which could determine whether or not a given proof was motivated. In any case, recognition of motivated proofs represents an easier sub-goal towards generating motivated proofs.
@@
For some proofs where there were incorrect but largely inconsequential components, we manually altered some equations to enforce correctness and removed any opening or closing sentences that included the phrase “motivated proof” to avoid misleading the model. As with generation, we prompted the model by explaining the definition of a motivated proof and giving an example and non-example for each of the other problems, where the order of examples and non-examples was alternated to ensure that the model wasn’t learning the ordering. Each model was given three attempts per proof per problem, and the results are recorded below. We use HM, HU, MM and MU as abbreviations for “human motivated”, “human unmotivated” and “machine unmotivated” respectively.
@@
Under this setup, o1-preview performed strongly, while Llama was heavily biased towards accepting a proof as motivated. A possible explanation for Llama’s poor performance in the MU set is that the machine-generated unmotivated proofs, unlike the human-generated ones, were attempts at generating motivated proofs, so these proofs contained some superficial features such as proof length and “chatty” language that the model had associated with motivated proofs.
@@
5 Mathematical Datasheet for Datasets
We next propose an amendment to the datasheets for datasets (DfD) framework  (Gebru et al., 2021), specific for datasets in mathematics. The DfD framework provides a comprehensive questionnaire that any dataset developer would benefit from answering; we encourage developers of new mathematical datasets to include all original DfD questions as well as our new set of questions outlined below, yielding mathematical datasheets for datasets (mDfDs). (We note that there have been other important calls for data labelling, e.g., data cards (Pushkarna et al., 2022) and data statements (Bender and Friedman, 2018); we choose DfD as a starting point given its comprehensiveness and generality.)
Based on the survey of the landscape of mathematical datasets provided in this article, we propose the inclusion of the additional questions for any dataset involving mathematics, as outlined in Figure 1. We include two additional sets of questions to consider on a case-by-case basis, depending on whether the dataset includes formal or informal mathematics. If both, we encourage providing information for each set of questions. We strive for only light—but we argue critical—additions to the DfD framework to avoid excess burden on research teams.
Questions for all mathematics datasets: Does your dataset cover only end results (statements and proofs), or (also) intermediate proof representations (which ones)? What is the precise focus of the dataset (proofs, workflows of various kinds, reasoning arguments, heuristics, etc.)? If proofs are included, are they motivated? Who confirmed that they are motivated? What is the type (degree of abstraction vs. proof complexity) and level of difficulty (e.g., grade school, high school, undergraduate, IMO, a mix) of the mathematical items? Who determined the difficulty rating? What scale did you use to determine difficulty rating (e.g., 5-point Likert scale)? Wich MSC codes represent your dataset best? (List all applicable ones) (Case-based) For datasets that include formal mathematics: Which formal mathematics language and corresponding ATP/ITP was used (e.g., Lean, Isabelle, Coq, Prover9, Vampire)? What version of it was used? (Ideally reference the specific commit – if available, such as in the case of Lean) Which dependencies are used? (That is, does the data set stand on its own, or does it depend on existing libraries of formal mathematics (e.g., Isabelle AFP or Mathlib)? If so, what version of the library are you using, and when was it last updated?) (Case-based) For datasets that include informal mathematics: Which items (statements, proofs, workflow steps, etc.) are checked for correctness? For proofs: On what level of detail are your datapoints? (The level of detail can vary dramatically, from a high-level sketch of a proof, down to a “formalizable” proof that has all details so that its formalization does not need any filling of mathematical gaps.) What procedure is used to check for correctness? Who (and how many people) checked correctness? If any other annotations are made about the informal mathematics that are not covered by the previous questions (e.g., whether a proof is aesthetically-pleasing, whether a specific workflow representation was used), please provide information on how (and by whom) they were procured. What language is the informal mathematics written in (e.g., English, German, French, Japanese, Chinese)?
Figure 1: The questionnaire for mDfDs. We note that some of the questions are dependent on each other, so all need to be answered. E.g., for the first question, what counts as an “undergraduate” differs from country to country (and even from university to university). The second question is thus necessary to put the first in context.
@@ 6 Conclusion
Regarding purely proof creation, time will tell whether human-annotated proofs are necessary–or whether a purely formal approach together with ingenious proof search techniques will succeed. The comparatively short history of machine learning has taught us to expect the unexpected. In the short term, however, it is imperative to establish better datasets and novel benchmarks for mathematics (either in natural language or formal language), to go beyond the current ones that only test proof or result creation - and also test the ability to assess different mathematical workflows, to summarize mathematics, to explain limitations of proof techniques etc. This will support the next generation of machine learning models and AI tools that can help us discover more mathematics fast and, subsequently assist with any other scientific discipline that uses mathematics as a foundation (e.g. systems biology that relies on ordinary differential equations, or physics that relies on several subfields of mathematics – and inspires new subfields at the same time). At the same time, better mathematical copilots may have strong educational benefits.
@@ Appendix A Appendix: Problems for Motivated Proof Experiments
For the motivated proof experiments in 4.3, we used five problems including the two examples. These are presented how they were presented to the models, including the human-written motivated and unmotivated proof used for in-context examples.
@@
MOTIVATED PROOF: Let $f:X\rightarrow\mathbb{P}(X)$ be our function, we need to find some $S\subseteq X$ not in the image of $f$. Not knowing which subset of $X$ to take, we can treat the subset as an unknown, just as we do when solving an equation, and try to narrow down the possibilities. The most general subset of $X$ can be expressed as $\{x\in X:P(x)\}$ for some as yet unspecified property $P$. We now want to prove, for an arbitrary element $y$ of $X$, that $f(y)\neq\{x\in X:P(x)\}$. To obtain our contradiction, we need either an element $x$ of $f(y)$ such that $\neg P(x)$ or an element $x$ of the complement of $f(y)$ such that $P(x)$. There are not many elements around, so trying $y$ is one of the first things to do, and then we find that we need either $y\in f(y)$ and $\neg P(y)$ or $y\notin f(y)$ and $P(y)$. The choice $P(x)=x\notin f(x)$ satisfies this, so the set $S=\{x\in X:x\notin X\}$ has the desired property.
@@
There exists a subset A of the natural numbers with cardinality n, such that the set $A+A={a_{1}+a_{2}:a_{i},a_{j}\in A}$ has cardinality $12n-$ 1.
@@
Trying to distinguish as many elements as possible, we fix some $i$ and take this inequality to its logical conclusion, finding that $a_{1}+a_{i}<....<a_{n}+a_{i}$, or that we have n distinct elements. However, we can further extend this chain to the left if $1<i$ or the right if $i<n$, giving us the longer chain $a_{1}+a_{1}<...<a_{1}+a_{i}<...<a_{n}+a_{1}<...<a_{n}+a_{n}$, which has $12n-$ 1 distinct elements. If we want to have $|A+A|=2n-1$, then these are all the elements of $A+A$, but we still have this free parameter $i$. Repeating the construction with some $j>i$ and observing the first syntactic different element, we notice that $a_{1}+a_{i+1}=a_{2}+a_{i}$, and isolating the variable i we notice that $a_{2}-a_{1}=a_{i+1}-a_{i}$. This defines an arithmetic progression with first element $a_{1}$ and common difference $a_{2}-a_{1}$, so let’s see if that enough.
Suppose A is an arithmetic progression of length $n$, then $A={a,a+d,\ldots,a+(n-1)d}$, then $A+A={2a,2a+d,\ldots,2a+(2n-2)d}$ has size $12n-$ 1, completing the proof.
UNMOTIVATED PROOF: Let $A=\{1,...,n\}$, which has cardinality n, then $A+A=\{2,...,2n\}$ has cardinality $12n-$ 1, completing the proof.
@@
But we also know that $x$ is nilpotent, so if we have $r$ such that $x^{r}=0$, then all coefficients from $x^{r}$ onwards can be ignored. This solves our issue, as we can let $m=r-1$ to remove the $a_{m}=0$ condition, and we are left with $a_{k}=(-1)^{k}$, so $(1+x)$ is a unit with inverse $\sum_{k=0}^{r-1}(-1)^{k}\cdot x^{k}$.
@@
MOTIVATED PROOF: We first try to simplify $\phi(n)/n$. Expressing $n$ in its prime factorisation to match the definition of $\phi(n)$, we have $\phi(n)/n=\prod_{i=1}^{m}(p_{i}^{k_{i}-1}(p_{i}-1)/p_{i}^{k_{i}})=\prod_{i=1}^{m}(1-1/p_{i})$. We notice that each factor is between 00 and $1$, so the product is large where there are few factors so we let $m=1$ (or equivalently, letting $n=p^{k}$), giving us $\phi(n)/n=(1-1/p)$. Since this is clearly less than $1$, we only need to choose $p$ such that $1-\epsilon<1-1/p$, which simplifies easily to $p>1/\epsilon$. Since there are infinitely many prime numbers, we can always choose such a $p$.
UNMOTIVATED PROOF: Let $p$ be the smallest prime number such that $p>1/\epsilon$, which exists as there are infinitely many primes. Then we have $\phi(p)/p=1-1/p$, and as $p>1/\epsilon$, we see that $1-\epsilon<1-1/p<1-\epsilon$ as required.

### v2

@@ 4.4 Motivated Proofs
We do not doubt that there is substantial value in bespoke AI tools that can automate specific elements of a mathematician’s workflow. However, if the goal is not only to provide correct proofs but also to enhance a user’s mathematical understanding of the results (zhang2023cogsciMath) and guide new discoveries, we need new tooling – and we argue, new kinds of data. In particular, we argue for data which faithfully represents the process of proof discovery instead of proof exposition. Towards this goal, we introduce motivated proofs (polya1949or; morris2019motivated), which contain and make transparent more of the proof discovery process. We argue for their value as a standard for LLM evaluation. We then provide some examples and discuss some preliminary observations about current LLM’s ability to construct and identify motivated proofs. Finally, we speculate on a path towards large-scale evaluation of a model’s ability to produce motivated proofs.
@@ 4.4.1 Examples of motivated proofs
But we also know that $x$ is nilpotent, so if we have $r$ such that $x^{r}=0$, then all coefficients from $x^{r}$ onwards can be ignored. This solves our issue, as we can let m=r1m=r-1 to remove the $a_{m}=0$ condition, and we are left with $a_{k}=(-1)^{k}$, so $(1+x)$ is a unit with inverse $\sum_{k=0}^{r-1}(-1)^{k}\cdot x^{k}$.
@@ 4.4.2 Motivated proofs as an evaluation metric
As argued in (polya1949or; morris2019motivated), a motivated proof is more informative to the reader, as it provides more insight into how the proof was discovered and how the reader might discover similar results themselves. More speculatively, a model capable of producing motivated proofs will likely generalize better to novel problems. The first reason for this conjecture is that most results in the literature are not fully motivated, and a model that is capable of recognising this is less likely to blindly recall its training data. Requiring proofs to be motivated also restricts the amount of brute-force search a model can perform, so performance depends more on reasoning ability and less on test-time computational power. A dedicated study to test whether models that generate motivated proofs indeed generalize better to novel problems could be worth pursuing, because it might reshape how we approach automated reasoning.
@@
In most responses, the model showed some evidence of “understanding" the concept of motivated proofs by avoiding unmotivated constructions. However, this often led to pseudo-motivations, where some justification was given before a construction, but it didn’t provide any useful substance besides obscuring the unmotivated step (and often appealed to some irrelevant fallacies). In the example below, the unmotivated step of Cantor’s theorem is relegated to the statement “An important question arises: Does $x$ belong to its own image f(x)f(x)?".
@@
Overall, o1-preview seemed to outperform Llama on most problems. We also observed a slight improvement by using negative as well as positive examples to “explain" the concept. As an unfortunate side effect, the extra information also increased the tendency to “obscure" unmotivated arguments as discussed above.
@@
To conduct a larger scale investigation of a model’s ability to generate motivated proofs, it would help if there was an existing “judge" model which could determine whether or not a given proof was motivated. In any case, recognition of motivated proofs represents an easier sub-goal towards generating motivated proofs.
@@
For some proofs where there were incorrect but largely inconsequential components, we manually altered some equations to enforce correctness and removed any opening or closing sentences that included the phrase “motivated proof" to avoid misleading the model. As with generation, we prompted the model by explaining the definition of a motivated proof and giving an example and non-example for each of the other problems, where the order of examples and non-examples was alternated to ensure that the model wasn’t learning the ordering. Each model was given three attempts per proof per problem, and the results are recorded below. We use HM, HU, MM and MU as abbreviations for “human motivated”, “human unmotivated” and “machine unmotivated” respectively.
@@
Under this setup, o1-preview performed strongly, while Llama was heavily biased towards accepting a proof as motivated. A possible explanation for Llama’s poor performance in the MU set is that the machine-generated unmotivated proofs, unlike the human-generated ones, were attempts at generating motivated proofs, so these proofs contained some superficial features such as proof length and “chatty" language that the model had associated with motivated proofs.
@@
Replication with newer models
Given the significant improvements in model performance since the release of an earlier preprint, we have replicated these investigations with newer models. Although these are still preliminary investigations, we acknowledge the risk of data contamination.
We first replicated the generation task, prompting each model to generate a motivated proof of each result given an example and non-example of each other result. Each task was repeated three times, with each response manually classified by the authors and the results tabulated below:
OpenAI
GPT-5.2 Pro
DeepSeek
Prover-V2
Google
Gemini 3 Pro
OpenAI
GPT-OSS 120B
Total
Cantor’s theorem 3/3 0/3 3/3 0/3 6/12
Small doubling 1/3 0/3 2/3 1/3 4/12
Integer sum 3/3 0/3 3/3 3/3 9/12
Nilpotent inverses 3/3 0/3 1/3 3/3 7/12
Large totients 3/3 3/3 3/3 3/3 12/12
Total 13/15 3/15 12/15 10/15 38/60
This simple investigation was sufficient to observe substantial variation between models, although we believe that more difficult problems are needed to meaningfully distinguish the top models. It is also noteworthy that both closed-source models consistently succeeded on Cantor’s Theorem, which neither of the older models could succeed on. It is also interesting that DeepSeek Prover-v2 performed poorly despite its mathematical specialization, which is consistent with the hypothesis that heavy fine-tuning for formal proof generation may reduce performance on stylistically different proofs.
We then repeated the binary classification task with newer models, concentrating on the more difficult and relevant machine-machine distinction, using proofs generated in the previous task. The first (motivated,unmotivated) pair was generated by GPT-5.2 Pro and DeepSeek Prover respectively, and the differences were stark enough that every model was correct 100% of the time. For the second pair we compared Gemini’s correct responses to a varied selection of incorrect model responses (including Gemini’s where possible). We used a human-written unmotivated proof for the problem of “large totients", since none of the newer models produced an unmotivated proof. The performance of each model on the second pair is shown below.
Theorem
OpenAI
GPT-5.2 Pro
DeepSeek
Prover-V2
Google
Gemini 3 Pro
OpenAI
GPT-OSS 120B
Total
Cantor’s theorem 2/3 1/3 3/3 1/3 7/12
Small doubling 3/3 0/3 0/3 3/3 6/12
Integer sum 3/3 3/3 3/3 3/3 12/12
Nilpotent inverses 3/3 1/3 3/3 3/3 10/12
Large totients 3/3 3/3 3/3 3/3 12/12
Total 14/15 8/15 12/15 13/15 47/60
This generally shows some improvement in distinguishing motivated from unmotivated proofs, although some more subtle distinctions are still consistently missed.
@@ 5 Conclusion
Regarding purely proof creation, time will tell whether human-annotated proofs are necessary–or whether a purely formal approach together with ingenious proof search techniques will succeed. The comparatively short history of machine learning has taught us to expect the unexpected. In the short term, however, it is imperative to establish better datasets and novel benchmarks for mathematics (either in natural language or formal language), to go beyond the current ones that only test proof or result creation - and also test the ability to assess different mathematical workflows, to summarize mathematics, to explain limitations of proof techniques etc. This will support the next generation of machine learning models and AI tools that can help us discover more mathematics fast and, subsequently assist with any other scientific discipline that uses mathematics as a foundation (e.g. systems biology that relies on ordinary differential equations, or physics that relies on several subfields of mathematics – and inspires new subfields at the same time). At the same time, better mathematical copilots may have strong educational benefits. Recent controlled experiments suggest that LLM-based explanations can positively impact learning for high-school level mathematics, particularly when students attempt problems independently before consulting the model (Kumar2025MathEducationLLMs). Moreover, Human-AI tutoring systems such as Tutor CoPilot have demonstrated measurable improvements in student outcomes when LLMs provide real-time pedagogical guidance to novice tutors (wang2024tutor).
@@ Appendix A Appendix: Problems for Motivated Proof Experiments
For the motivated proof experiments in 4.4, we used five problems including the two examples. These are presented how they were presented to the models, including the human-written motivated and unmotivated proof used for in-context examples.
@@
MOTIVATED PROOF: Let $f:X\rightarrow\mathbb{P}(X)$ be our function, we need to find some $S\subseteq X$ not in the image of $f$. Not knowing which subset of $X$ to take, we can treat the subset as an unknown, just as we do when solving an equation, and try to narrow down the possibilities. The most general subset of $X$ can be expressed as $\{x\in X:P(x)\}$ for some as yet unspecified property $P$. We now want to prove, for an arbitrary element $y$ of $X$, that $f(y)\neq\{x\in X:P(x)\}$. To obtain our contradiction, we need either an element $x$ of $f(y)$ such that $\neg P(x)$ or an element $x$ of the complement of $f(y)$ such that $P(x)$. There are not many elements around, so trying $y$ is one of the first things to do, and then we find that we need either $y\in f(y)$ and $\neg P(y)$ or $y\notin f(y)$ and $P(y)$. The choice $P(x)=x\notin f(x)$ satisfies this, so the set $S=\{x\in X:x\notin f(x)\}$ has the desired property.
@@
There exists a subset A of the natural numbers with cardinality n, such that the set $A+A={a_{1}+a_{2}:a_{i},a_{j}\in A}$ has cardinality 2n12n-1.
@@
Trying to distinguish as many elements as possible, we fix some $i$ and take this inequality to its logical conclusion, finding that $a_{1}+a_{i}<....<a_{n}+a_{i}$, or that we have n distinct elements. However, we can further extend this chain to the left if $1<i$ or the right if $i<n$, giving us the longer chain $a_{1}+a_{1}<...<a_{1}+a_{i}<...<a_{n}+a_{1}<...<a_{n}+a_{n}$, which has 2n12n-1 distinct elements. If we want to have |A+A|=2n1|A+A|=2n-1, then these are all the elements of $A+A$, but we still have this free parameter $i$. Repeating the construction with some $j>i$ and observing the first syntactic different element, we notice that $a_{1}+a_{i+1}=a_{2}+a_{i}$, and isolating the variable i we notice that $a_{2}-a_{1}=a_{i+1}-a_{i}$. This defines an arithmetic progression with first element $a_{1}$ and common difference $a_{2}-a_{1}$, so let’s see if that enough.
Suppose A is an arithmetic progression of length $n$, then $A={a,a+d,\ldots,a+(n-1)d}$, then $A+A={2a,2a+d,\ldots,2a+(2n-2)d}$ has size 2n12n-1, completing the proof.
UNMOTIVATED PROOF: Let $A=\{1,...,n\}$, which has cardinality n, then $A+A=\{2,...,2n\}$ has cardinality 2n12n-1, completing the proof.
@@
But we also know that $x$ is nilpotent, so if we have $r$ such that $x^{r}=0$, then all coefficients from $x^{r}$ onwards can be ignored. This solves our issue, as we can let m=r1m=r-1 to remove the $a_{m}=0$ condition, and we are left with $a_{k}=(-1)^{k}$, so $(1+x)$ is a unit with inverse $\sum_{k=0}^{r-1}(-1)^{k}\cdot x^{k}$.
@@
MOTIVATED PROOF: We first try to simplify $\phi(n)/n$. Expressing $n$ in its prime factorisation to match the definition of $\phi(n)$, we have $\phi(n)/n=\prod_{i=1}^{m}(p_{i}^{k_{i}-1}(p_{i}-1)/p_{i}^{k_{i}})=\prod_{i=1}^{m}(1-1/p_{i})$. We notice that each factor is between 0 and $1$, so the product is large where there are few factors so we let $m=1$ (or equivalently, letting $n=p^{k}$), giving us $\phi(n)/n=(1-1/p)$. Since this is clearly less than $1$, we only need to choose $p$ such that $1-\epsilon<1-1/p$, which simplifies easily to $p>1/\epsilon$. Since there are infinitely many prime numbers, we can always choose such a $p$.
UNMOTIVATED PROOF: Let $p$ be the smallest prime number such that $p>1/\epsilon$, which exists as there are infinitely many primes. Then we have $\phi(p)/p=1-1/p$, and as $p>1/\epsilon$, we see that $1-\epsilon<1-1/p<1$ as required.